OpenStudy (anonymous):

how tp write 7 ^1/2 in radical form

OpenStudy (anonymous):

sqrt(7)

OpenStudy (ybarrap):

$$\sqrt{7}$$

OpenStudy (amistre64):

\[a^{1/n}=\sqrt[n]{a}\]